OpenStudy (anonymous):

If the width of a rectangular piece of carpet is 2x + 1 yards and the length is three yards less than twice as long as the width, what is the area of the carpet? A. 6x² + 2x – 1 B. 8x² + 2x – 1 C. 8x² – 2x – 1 D. 8x² + 6x – 1

OpenStudy (anonymous):

Width=2x+1 Length=(2*Width)-3 =2(2x+1)-3=4x+2-3=4x-1 Now can u find area (Area=Width*Length)